In a supermarket,There are 30 cartons of eggs on a shelf. the number of cracked eggs in each carton is shown in the following table
Number of cracked eggs 0 1 2 3
Number of cartons 21 6 2 1
If a carton of eggs is selected at random,find the probabilities of the following events happening:
The carton contains eggs without any cracks

Answer by ewatrrr(24785)


Hi

Eggs Broken (x)	0	1	2	3
30 cartons	21	6	2	1
 P(x)          21/30   6/30   2/30	1/30 and sum = 30/30 = 1

P(0 cracked) = 21/30 = 7/10 = .70  0r 70%

Wish You the Best in your Studies.